OpenStudy (anonymous):

Polive agents flying a constant 200km/h horizontally in a low flying airplane wish to drop an explosive onto a master criminal's car travelling 130km/h on a level highway 88.0m below. At what angle (with the horizontal) should the car be in their sights when the bomb is released?

OpenStudy (anonymous):



OpenStudy (anonymous):

h=88 m No matter when the bomb is dropped it will strike the ground in time ,t \[t = \sqrt{2h/g}\] Suppose the bomb is dropped when the relative separation between the bomb and the car is x. The velocity of bomb as seen by the car will be 200-130 =70km/h and it has to cover x distance in the above time with this speed. So x = v*t Make sure that you convert speed in m/s. Then you will get x in meters. You know x and h and you can easily find the angle.